OpenStudy (anonymous):

MEDAL AND FAN!!!! Find the sum: -12 + (-15) = A. -27 B. -3 C. 3 D. 27

OpenStudy (anonymous):

-12+(-15)=-12-15=-27 Answer is A

OpenStudy (anonymous):

THANKS